In a particular year my sister was twice as old as I was and I was twice as old as my brother. The product of our ages was 216. The sum of our ages must have been?
a. 21 b. 26 c. 34 d. 54

Answer by ankor@dixie-net.com(22740)

let h = his age in a particular yr
let s = sister's age
let b = brother's age
:
Write an equation for each statement
:
In a particular year my sister was twice as old as I was
s = 2h
and I was twice as old as my brother.
h = 2b
or (get s and b in terms of h)
b = .5h
The product of our ages was 216.
h*s*b = 216
replace s and b
h * 2h * .5h = 216
h^3 = 216
find the cube root
h = 6 yrs old
then
b = 3 yrs old
and
s = 12 yrs old
:
6 + 3 + 12 = 21